Thin films morrning report 5/11/04

james haydon jhaydon at
Tue May 11 09:33:55 PDT 2004

Thinfilms Morning Report 5/11/2004



Gryphon:            fixed power supplies installed and seams to be working

Innotec;            ok

Balzers :            ok

STS Dep:            ok

Implanter:            ok

Metalica:            new cryo pump installed and working at least as 
well as the old one

SCT:                completion of installation near end of May

CPD:                ok                   

WbGen-CTB:            new control temperature bath installed

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <>

More information about the thinfilms mailing list